OpenStudy (anonymous):

Suppose F(x)=g(h(x)) . If g(2)=3 , g′(2)=2 , h(0)=2 , and h′(0)=6 find F′(0) .

OpenStudy (goformit100):

just diffrentiate thee function

OpenStudy (anonymous):

similar to the last one \[F(x)=g(h(x))\] \[F'(x)=g'(h(x))\times h'(x)\] by the chain rule

OpenStudy (anonymous):

\[F'(0)=g'(h(0))\times h'(0)\] fill in the numbers you know to get the answer you want

OpenStudy (anonymous):

2*2*6=24 but it says my answer is incorrect